Fermi National Laboratory
 

Operation Summary for April 15, 2011 - April 18, 2011

 

The Integrated Luminosity for the period from 4/11/11 to 4/18/11 was 36.6 Pb-1 (inverse picobarns). NuMI took 4.9E18 of beam during this same time period.

Friday April 15

The day shift began with the Tevatron (TEV) in a proton-only study store, with Operations monitoring the Recycler stash (311.7E10), with the Antiproton Source (Pbar) stacking (296.4E10), with the Main Injector (MI), Linac, and Booster operating normally, with the NuMI, MiniBooNE, and Meson FTFB T-978 experiments taking beam, and with the Collider Detector at Fermilab (CDF) and D-Zero (D0) waiting for the next store.

During the day shift, Operations halted stacking at 9:30 AM so that Pbar personnel could work on their dump LCW cooling system. The Pbar target blower tripped off; system personnel investigated and repaired it. Operations put the TEV into HEP shot setup and established store 8662 at 12:50 PM with an average initial luminosity (AIL) of 212.0E30. Pbar resumed stacking at 3:05 PM. A TEV RF station (TRF2) tripped off; RF personnel brought the station back online. Operations lowered the NuMI intensity to 1.4E13 protons per pulse.

During the evening shift, Operations disabled beam to T-978 at 6 PM. The Pbar target blower failed again; Pbar personnel repaired it and stacking resumed at 6:30 PM. Operations terminated store 8662 at 9:05 PM and established store 8663 at 10:20 PM with an AIL of 271.0E30. The Booster RF East Anode power supply tripped off and caused BRF4 to trip off; BRF4 would not reset. Operators brought BRF6 on in its place.

Saturday April 16

The midnight shift began with Operations monitoring the store 8663 (187.91E30) and the stash (63.1E10), with Pbar stacking (17.9E10), and with NuMI and MiniBooNE taking beam.

During the midnight shift, Operations established beam to T-978 at 5:45 AM. Operations reported that MiniBooNE pulses attached to the Switchyard 120 GeV event were causing beam position problems at the MI-8 line collimators that were in turn causing MiniBooNE and Main Injector beam permit trips. Operations decreased the MiniBooNE beam intensity from 12 Booster turns to 11 Booster turns.

During the day shift, Operations terminated store 8663 at 2:35 PM, put the TEV into shot setup, and established store 8664 at 4 PM with an AIL of 310.0E30.

During the evening shift, Operations disabled beam to T-978 at 6 PM.

Sunday April 2017

The midnight shift began with Operations monitoring store 8664 and the stash, with Pbar stacking, and with NuMI and MiniBooNE taking beam.

During the midnight shift, Operations established beam to T-978 at 4:40 AM.

During the day shift, Operations terminated store 8664 at 8:30 AM, put the TEV into shot setup, and established store 8665 at 9:50 AM with an AIL of 325.0E30.

During the evening shift, Operations monitored the accelerator complex.

Monday April 18

The midnight shift began with Operations monitoring store 8665 and the stash, with Pbar stacking, and with NuMI and MiniBooNE taking beam.

During the midnight shift, Operations terminated store 8665 at 3:15 AM, put the TEV into shot setup, and established store 8666 at 4:43 AM with an AIL of 333.0E30. Operations established beam to T-978 at 5:50 AM.

The Plans for Monday

The plans for today are to stack, stash, and store. The Run Coordinator will allow Booster personnel to repair BRF4 and bring it back on line and to conduct a study. He will also allow MI personnel to conduct a hipot.

 

Accelerator Update Archive

More Information
For Tevatron luminosity charts and the current status of Fermilab's accelerators and detectors (live!), please go to Fermilab Now.

Comments and Suggestions
What do you think about the Accelerator Updates? Please send comments and suggestions to: accelupdates@fnal.gov.

last modified 04/18/2011    email Fermilab
Security, Privacy, LegalFermi National Accelerator Laboratory